# Dimensions and metric dyadic cubes

### v1

@@ 1. Introduction
All three notions of dimension have proved to be extremely useful, and are often studied simultaneously in situations where they differ. However, their original definitions can be quite complicated for the sake of calculations. The main idea behind the Hausdorff dimension $\dim_{H}E$ of a set $E\subset\mathbb{R}^{n}$ is that the dimension of the set is essentially the critical exponent $s\geq 0$ for which the $s$-dimensional Hausdorff measure of $E$, denoted by $H^{s}(E)$, changes from being infinite to being 00. This is motivated by the property of the Lebesgue measure that, for instance, assigns infinite length and zero volume to a $2$-dimensional object. On the other hand, the Minkowski dimension $\dim_{B}E$ of a bounded set $E\subset\mathbb{R}^{n}$ is intuitively the exponent $s>0$ for which $N(E,r)$, i.e. the minimum number of sets of diameter at most $r$ needed to cover $E$, is approximately $r^{-s}$, for all small $r>0$. The Assouad dimension $\dim_{A}E$ of $E$ is defined by utilizing the idea of the Minkowski dimension after “zooming in” (or “zooming out”) at all points of $E$. For $\theta\in(0,1)$, the $\theta$-Assouad spectrum $\dim_{A}^{\theta}E$ of $E$ is defined similarly to the Assouad dimension, with a restriction on the allowed zoom scales that depends on $\theta$. We refer to Section 2 for rigorous definitions.
In the Euclidean setting, one feature of $\mathbb{R}^{n}$ that has facilitated the study of dimensions, especially within applied areas, has been the simplification of the aforementioned definitions using dyadic cubes. In particular, the Hausdorff dimension of $E$ is equal to the critical exponent $s\geq 0$ for which the $s$-dimensional cubic measure $M^{s}(E)$ changes from being infinite to being equal to 00. The cubic measure $M^{s}(E)$ is defined using dyadic cubes, unlike the Hausdorff measure which uses arbitrary sets. Similarly, for the definition of the Minkowski and Assouad dimension, if $E$ is contained in a ball $B(x,R)$ and its circumscribed cube $Q(x,R)$, one can replace the number $N(E,r)$ by the smallest number of dyadic sub-cubes of $Q(x,R)$ of level $m$ needed to cover $E$, denoted by $D(E,m)$, and ask that $D(E,m)\simeq 2^{ms}$, for all large $m\in\mathbb{N}$. Such simplifications have been an invaluable tool in many applications of fractal geometry. In particular, in the area of dimension distortion under mapping classes, the simplified definitions have been used extensively (for instance [20, 26, 22]), including in previous work of the author in [8, 7], and jointly by Tyson and the author in [11, 10].
@@
Let $(X,d)$ be a doubling metric space. There is a finite collection of dyadic cube systems $\mathcal{S}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$ of $X$ such that the following hold:
@@
    There is a constant $\widetilde{C}>0$ such that for any $\theta\in(0,1)$, the Assouad spectrum $\dim_{A}^{\theta}E$ is equal to the infimum of all $\alpha>0$ for which there is $C>0$ with
@@
    for all $x\in E$, $R>0,m\in\mathbb{N}$ with $10<\widetilde{C}\delta^{m}R\leq R^{1/\theta}<R<$ 1.
@@
    There is a constant $\widetilde{C}>0$ such that the Assouad dimension $\dim_{A}E$ is equal to the infimum of all $\alpha>0$ for which there is $C>0$ with
@@
    for all $x\in E$, $R>0$, $m\in\mathbb{N}$ with $\widetilde{C}\delta^{m}R\leq R$.
@@ 2. Background
Let $(X,d)$ be a metric space. We use the Polish notation $d(x,y)=|x-y|$ for all $x,y\in X$ and denote the open ball centered at $x$ of radius $r>0$ by $B(x,r):=\{z\in X:\,\,|x-z|<r\}.$ Given a ball $B=B(x,r)\subset X$, we denote by $\lambda B$ the ball $B(x,\lambda r)$, for $\lambda>0$. Given a subset $U$ of $X$, we denote by $|U|$ the diameter of $U$.
@@
$\dim_{A}E=\inf\left\{\alpha>0\,:\,{\exists\,C>0\mbox{ s.t. }N(E\cap B(x,R),r)\leq C(R/r)^{\alpha}\atop\mbox{ for all $0<r\leq R$ and all $x\in E$}}\right\}.$
@@
with the inequalities being strict in various cases (see for instance [6]). The potential “gap” between the Minkowski and the Assouad dimension received increased attention by Fraser and Yu [19], who defined a collection of dimension notions that reside in that gap. For $10<\theta<$ 1 and a set $E\subset X$, define
$\dim_{A}^{\theta}E=\inf\left\{\alpha>0\,:\,{\exists\,C>0\mbox{ s.t. }N(B(x,R)\cap E,r)\leq C(R/r)^{\alpha}\atop\mbox{ for all $0<r\leq R^{1/\theta}<R<1$ and all $x\in E$}}\right\}.$
@@
In the context of the typical Euclidean metric spaces $\mathbb{R}^{n}$, it is quite elementary to equivalently express the dimensions $\dim_{H}E$ and $\dim_{B}E$ using dyadic cubes instead of arbitrary sets, by modifying the quantities $H^{s}_{r}(E)$ and $N(E,r)$. This idea was extended by Tyson and the author in [11] for the Assouad dimension and the Assouad spectrum. In this subsection we present and define the appropriate dyadic cube notion in the context of metric spaces, in order to prove similar expressions for the aforementioned dimensions.
@@
Suppose $(X,d)$ is a doubling metric space. Let $0<c_{0}\leq C_{0}<\infty$ and $\delta\in(0,1)$ with $12C_{0}\delta\leq c_{0}$. For any non-negative $k\in\mathbb{Z}$ and collection of points $\{z_{i}^{k}\}_{i\in I_{k}}$ with
@@
we can construct a collection of sets $\{Q_{i}^{k}\}_{i\in I_{k}}$ such that
@@
Fix a doubling metric space $(X,d)$ with doubling constant $C_{d}$, and constants $\delta<1/100$, $c_{0}$ and $C_{0}$, as in Theorem A, for the rest of the paper. Moreover, for every non-negative $k\in\mathbb{Z}$ we fix a collection of points $\{z_{i}^{k}\}_{i\in I_{k}}$. To see why such a collection of points exists, consider the covering $\{B(z,c_{0}\delta^{k}):z\in X\}$ of $X$ and apply the $5B$-covering lemma (see [21]). By separability of $X$ and by choosing $c_{0}$ and $C_{0}$ so that $5c_{0}\delta^{k}<C_{0}\delta^{k}$, the existence of centers $\{z_{i}^{k}\}_{i\in I_{k}}$ is ensured.
By [24, Theorem 4.1], we can fix finitely many dyadic cube systems of $X$ satisfying Theorem A, say $\mathcal{S}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$ with $K=K(C_{d},\delta)\in\mathbb{N}$, so that for every $x\in X$, $R>0$, there are a constant $C_{\delta}>0$ and a $\delta$-dyadic cube $Q(x,R)\in\mathcal{D}_{t}$, for some $t=t_{x,R}\in\{1,\dots,K\}$, so that $B(x,R)\subset Q(x,R)$ and
@@
We have introduced all the necessary notions to rigorously define the cubic measure and dyadic cube covering number in the setting of metric spaces, which are stated in Theorem 1.1. Given $x\in X$, $R>0$, $m\in\mathbb{N}$, denote by $W_{m}(Q(x,R))\subset\mathcal{D}_{t}$ the dyadic cubes of level $L_{R}+m$ that are contained in $Q(x,R)\in\mathcal{D}_{t}$, where $L_{R}$ is the level of the circumscribed cube $Q(x,R)$ of the ball $B(x,R)$. Note that by Theorem A (iii) and (2.3), the level $L_{R}$ of $Q(x,R)$ does not depend on the point $x$. Given a set $E\subset X$, we denote by $D(E\cap B(x,R),m)$ the smallest number of cubes in $W_{m}(Q(x,R))$ needed to cover $E\cap B(x,R)$. If $E\subset B(x,R)$, we simplify the notation and write $D(E,m)$.
@@ 3.1. Hausdorff dimension and cubic measure
Due to (2.1) and the one-to-one correspondence of cubes of level $m_{i}$ and centers $z_{\ell}^{m_{i}}$, along with $\dim_{A}X<\infty$ imply that there can be at most $C^{\prime}>0$ cubes of level $m_{i}$ intersecting the ball $B(x_{i},4C_{0}\delta^{m_{i}})$, for some constant $C^{\prime}$ that depends only on $\delta$ and the doubling constant $C_{d}$ of $X$. By (2.2), these cubes in fact cover the ball $B(x_{i},4C_{0}\delta^{m_{i}})$. An additional application of the doubling condition of $X$ yields that every dyadic cube of level $m_{i}$ intersecting $B(x_{i},4C_{0}\delta^{m_{i}})$ can be decomposed into at most $C^{\prime\prime}>0$ cubes of level $m_{i}+1$, with $C^{\prime\prime}$ depending only on $\delta$ and $C_{d}$. Denote this collection of cubes of level $m_{i}+1$ by $\{Q_{j}^{m_{i}+1}\}_{j}$. Note that by the right-hand side of (3.1) we have $U_{i}\subset B(x_{i},4C_{0}\delta^{m_{i}})$, for every $i$. Thus, since $\{U_{i}\}_{i}$ is a cover of $E$, we have
@@
with $|Q_{j}^{m_{i}+1}|\leq r$ by Theorem A (iii), (3.1), and choice of $\{U_{i}\}$. In addition, by (3.1) we have
@@ 3.2. Dyadic covering numbers and dimensions
$A_{\theta}=\left\{\alpha>0\,:\,{\exists\,C>0\mbox{ s.t. }N(B(x,R)\cap E,r)\leq C(R/r)^{\alpha}\atop\mbox{ for all $0<r\leq R^{1/\theta}<R<1$ and all $x\in E$}}\right\},$
@@
for which we need to show that $\dim_{A}^{\theta}E=\inf B_{\theta}$. We do so by proving that in fact $A_{\theta}=B_{\theta}$.
Let $x\in E$, $R>0$, $m\in\mathbb{N}$ and a cube $Q_{i}\in W_{m}(Q(x,R))\in\mathcal{D}$, for some $\mathcal{D}\in\mathcal{S}$ where the circumscribed cube $Q(x,R)$ is contained. By definition of $W_{m}(Q(x,R))$, the cube $Q_{i}$ is of level $L_{R}+m$, where $L_{R}$ is the level of $Q(x,R)$. Since $|Q(x,R)|\geq 3^{-1}c_{0}\delta^{L_{R}}$ by Theorem A (iii), we have by (2.3) and another application of Theorem A (iii) on $Q_{i}$ that
@@ 4. Final Remarks
The construction of dyadic cube systems of Hytönen-Kairema in [24] is actually given for quasimetric spaces. As a result, Theorem 1.1 is also true if $X$ is a quasimetric doubling space. The proofs are almost identical, with the only difference being the dependence of a few of the uniform constants on the quasimetric constant of the space.

### v2

@@ 1. Introduction
All three notions of dimension have proved to be extremely useful, and are often studied simultaneously in situations where they differ. However, their original definitions can be quite complicated for the sake of calculations. The main idea behind the Hausdorff dimension $\dim_{H}E$ of a set $E\subset\mathbb{R}^{n}$ is that the dimension of the set is essentially the critical exponent $s\geq 0$ for which the $s$-dimensional Hausdorff measure of $E$, denoted by $H^{s}(E)$, changes from being infinite to being 0. This is motivated by the property of the Lebesgue measure that, for instance, assigns infinite length and zero volume to a $2$-dimensional object. On the other hand, the Minkowski dimension $\dim_{B}E$ of a bounded set $E\subset\mathbb{R}^{n}$ is intuitively the exponent $s>0$ for which $N(E,r)$, i.e. the minimum number of sets of diameter at most $r$ needed to cover $E$, is approximately $r^{-s}$, for all small $r>0$. The Assouad dimension $\dim_{A}E$ of $E$ is defined by utilizing the idea of the Minkowski dimension after “zooming in” (or “zooming out”) at all points of $E$. For $\theta\in(0,1)$, the $\theta$-Assouad spectrum $\dim_{A}^{\theta}E$ of $E$ is defined similarly to the Assouad dimension, with a restriction on the allowed zoom scales that depends on $\theta$. We refer to Section 2 for rigorous definitions.
In the Euclidean setting, one feature of $\mathbb{R}^{n}$ that has facilitated the study of dimensions, especially within applied areas, has been the simplification of the aforementioned definitions using dyadic cubes. In particular, the Hausdorff dimension of $E$ is equal to the critical exponent $s\geq 0$ for which the $s$-dimensional cubic measure $M^{s}(E)$ changes from being infinite to being equal to 0. The cubic measure $M^{s}(E)$ is defined using dyadic cubes, unlike the Hausdorff measure which uses arbitrary sets. Similarly, for the definition of the Minkowski and Assouad dimension, if $E$ is contained in a ball $B(x,R)$ and its circumscribed cube $Q(x,R)$, one can replace the number $N(E,r)$ by the smallest number of dyadic sub-cubes of $Q(x,R)$ of level $m$ needed to cover $E$, denoted by $D(E,m)$, and ask that $D(E,m)\simeq 2^{ms}$, for all large $m\in\mathbb{N}$. Such simplifications have been an invaluable tool in many applications of fractal geometry. In particular, in the area of dimension distortion under mapping classes, the simplified definitions have been used extensively (for instance [20, 26, 22]), including in previous work of the author in [8, 7], and jointly by Tyson and the author in [11, 10].
@@
Let $(X,d)$ be a doubling metric space. There is a finite collection of $\delta$-dyadic cube systems $\mathcal{S}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$ of $X$, such that for every non-empty $E\subset X$, the following hold:
@@
    There is a constant $\widetilde{C}>0$, independent of $E$, such that for any $\theta\in(0,1)$, the Assouad spectrum $\dim_{A}^{\theta}E$ is equal to the infimum of all $\alpha>0$ for which there is $C>0$ with
@@
    for all $x\in E$, $R>0,m\in\mathbb{N}$ with $0<\widetilde{C}\delta^{m}R\leq R^{1/\theta}<R<1$, and $D(E\cap B(x,R),m)$ with respect to the circumscribed cube $Q(x,R)$.
@@
    There is a constant $\widetilde{C}>0$, independent of $E$, such that the Assouad dimension $\dim_{A}E$ is equal to the infimum of all $\alpha>0$ for which there is $C>0$ with
@@
    for all $x\in E$, $R>0$, $m\in\mathbb{N}$ with $\widetilde{C}\delta^{m}R\leq R$, and $D(E\cap B(x,R),m)$ with respect to the circumscribed cube $Q(x,R)$.
@@
Acknowledgements
I wish to thank Alex Rutar for the interesting conversations regarding this note. I am also grateful to the anonymous referee, whose comments improved the exposition of the note.
@@ 2. Background
Let $(X,d)$ be a metric space. We use the Polish notation |xy|:=d(x,y)|x-y|:=d(x,y) for all $x,y\in X$ and denote the open ball centered at $x$ of radius $r>0$ by $B(x,r):=\{z\in X:\,\,|x-z|<r\}.$ Given a ball $B=B(x,r)\subset X$, we denote by $\lambda B$ the ball $B(x,\lambda r)$, for $\lambda>0$. Given a subset $U$ of $X$, we denote by $|U|$ the diameter of $U$. While for every $x\in X$, $r>0$, we have $|B(x,r)|\leq 2r$, the radius of a ball in a metric space is not uniquely determined. However, for the rest of this manuscript we make the convention that for every ball $B$ that contains at least two distinct points, all radii of $B$ that we consider do not exceed $2|B|$. Thus, the notation $B(x,r)$ always implies that
(2.1) $2^{-1}|B(x,r)|\leq r\leq 2|B(x,r)|.$
@@
[…]
@@
with the inequalities being strict in various cases (see for instance [6]). The potential “gap” between the Minkowski and the Assouad dimension received increased attention by Fraser and Yu [19], who defined a collection of dimension notions that reside in that gap. For $0<\theta<1$ and a set $E\subset X$, define
@@
Thus, $\dim_{A}^{\theta}E$ is defined by the same process as $\dim_{A}E$, but with the restriction that the two scales $r$ and $R$ involved in the definition of the latter are related by the inequality $r^{\theta}\leq R$. The function $\theta\mapsto\dim_{A}^{\theta}E$ is called the (regularized) [^1] This definition can also be found as upper Assouad spectrum in the literature, while the original spectrum in [19] was defined with $r=R^{1/\theta}$. Due to the relation $r\leq R^{1/\theta}$ ensuring that the spectrum is monotone in $\theta$ for a fixed $E$, thus resulting in a more regular function of $\theta$, the term “regularized” was suggested by Tyson and the author. See also the discussion in [11]. Assouad spectrum of $E$. We often abuse the terminology and refer to a value $\dim_{A}^{\theta}E$ for some $\theta\in(0,1)$ as the Assouad spectrum of $E$.
@@
In the context of the typical Euclidean metric spaces $\mathbb{R}^{n}$, it is quite elementary to equivalently express the dimensions $\dim_{H}E$ and $\dim_{B}E$ using dyadic cubes instead of arbitrary sets, by modifying the quantities $H^{s}_{r}(E)$ and $N(E,r)$. This idea was also extended for the Assouad dimension and the Assouad spectrum (see [11, Proposition 2.5] for a proof). In this subsection we present and define the appropriate dyadic cube notion in the context of metric spaces, in order to prove similar expressions for the aforementioned dimensions.
@@
Suppose $(X,d)$ is a doubling metric space. Let $0<c_{0}\leq C_{0}<\infty$ and $\delta\in(0,1)$ with $12C_{0}\delta\leq c_{0}$. For any $k\in\mathbb{N}$ and collection of points $\{z_{i}^{k}\}_{i\in I_{k}}$ with
@@
where $I_{k}$ is an index set, we can construct a collection of sets $\{Q_{i}^{k}\}_{i\in I_{k}}$ such that
@@
Fix a doubling metric space $(X,d)$ with doubling constant $C_{d}$, and constants $\delta<1/100$, $c_{0}$ and $C_{0}$, as in Theorem A, for the rest of the paper. Moreover, for every $k\in\mathbb{N}$ we fix a collection of points $\{z_{i}^{k}\}_{i\in I_{k}}$ satisfying (2.2), (2.3). To see why such a collection of points exists, consider the covering $\{B(z,c_{0}\delta^{k}):z\in X\}$ of $X$ and apply the $5B$-covering lemma (see [21]). By separability of $X$ and by choosing $c_{0}$ and $C_{0}$ so that $5c_{0}\delta^{k}<C_{0}\delta^{k}$, the existence of centers $\{z_{i}^{k}\}_{i\in I_{k}}$ is ensured.
By [24, Theorem 4.1], we can fix finitely many dyadic cube systems of $X$ satisfying Theorem A, say $\mathcal{S}=\{\mathcal{D}_{1},\dots,\mathcal{D}_{K}\}$ with $K=K(C_{d},\delta)\in\mathbb{N}$, so that for every ball $B(x,R)$, $x\in X$, $R>0$, which contains at least two points, there are a constant $C_{\delta}>0$ and a $\delta$-dyadic cube $Q(x,R)\in\mathcal{D}_{t}$, for some $t=t_{x,R}\in\{1,\dots,K\}$, so that $B(x,R)\subset Q(x,R)$ and
@@
We have introduced all the necessary notions to rigorously define the cubic measure and dyadic cube covering number in the setting of metric spaces, which are stated in Theorem 1.1. We further assume without loss of generality that all balls considered in what follows contain at least two points, even if not explicitly stated, since all covering number conditions in the definitions of the Assouad dimension and spectrum are trivial otherwise. Given $x\in X$, $R>0$, $m\in\mathbb{N}$, denote by $W_{m}(Q(x,R))\subset\mathcal{D}_{t}$ the dyadic cubes of level $L_{R}+m$ that are contained in $Q(x,R)\in\mathcal{D}_{t}$, where $L_{R}$ is the level of the circumscribed cube $Q(x,R)$ of the ball $B(x,R)$. Note that by the convention $|B(x,R)|\geq R/2$, Theorem A (iii) and (2.4), the level $L_{R}$ of $Q(x,R)$ does not depend on the point $x$. Given a set $E\subset X$, we denote by $D(E\cap B(x,R),m)$ the smallest number of cubes in $W_{m}(Q(x,R))$ needed to cover $E\cap B(x,R)$. If $E\subset B(x,R)$, we simplify the notation and write $D(E,m)$.
@@ 3.1. Hausdorff dimension and cubic measure
Note that we assume that $|U_{i}|>0$, otherwise it would not contribute to the sum. Due to (2.2) and the one-to-one correspondence of cubes of level $m_{i}$ and centers $z_{\ell}^{m_{i}}$, along with $\dim_{A}X<\infty$ imply that there can be at most $C^{\prime}>0$ cubes of level $m_{i}$ intersecting the ball $B(x_{i},4C_{0}\delta^{m_{i}})$, for some constant $C^{\prime}$ that depends only on $\delta$ and the doubling constant $C_{d}$ of $X$. By (2.3), these cubes in fact cover the ball $B(x_{i},4C_{0}\delta^{m_{i}})$. An additional application of the doubling condition of $X$ yields that every dyadic cube of level $m_{i}$ intersecting $B(x_{i},4C_{0}\delta^{m_{i}})$ can be decomposed into at most $C^{\prime\prime}>0$ cubes of level $m_{i}+1$, with $C^{\prime\prime}$ depending only on $\delta$ and $C_{d}$. Denote this collection of cubes of level $m_{i}+1$ by $\{Q_{j}^{m_{i}+1}\}_{j}$. Note that by the right-hand side of (3.1) we have $U_{i}\subset B(x_{i},4C_{0}\delta^{m_{i}})$, for every $i$. Thus, since $\{U_{i}\}_{i}$ is a cover of $E$, we have
@@
with $|Q_{j}^{m_{i}+1}|\leq 4C_{0}\delta^{m_{i}+1}\leq r$ by Theorem A (iii), (3.1), and choice of $\{U_{i}\}$. In addition, by (3.1) we have
@@ 3.2. Dyadic covering numbers and dimensions
  2. (ii)
    It should be noted that in the proofs of Theorem 1.1 (i) and (ii), the fact that the dyadic system $\mathcal{D}$ is one of those lying in $\mathcal{S}$ is not used. Hence, the Hausdorff and Minkowski dimensions can in fact be expressed using any $\delta$-dyadic system with the properties listed in Theorem A.
@@
[…]
@@
for which we need to show that $\dim_{A}^{\theta}E=\inf B_{\theta}$. We do so by proving that in fact $A_{\theta}=B_{\theta}$. We also emphasize that even if we temporarily consider balls $B=B(x,R)$ in the definitions of the sets $A_{\theta},B_{\theta}$ against the convention (2.1), i.e. with $R>2|B(x,R)|$, then due to $N(B(x,R)\cap E,r)=N(B(x,2|B|)\cap E,r)$ the proof still reduces to considering only balls according to the convention.
Let $x\in E$, $R>0$, $m\in\mathbb{N}$ and a cube $Q_{i}\in W_{m}(Q(x,R))\in\mathcal{D}$, for some $\mathcal{D}\in\mathcal{S}$ where the circumscribed cube $Q(x,R)$ is contained. By definition of $W_{m}(Q(x,R))$, the cube $Q_{i}$ is of level $L_{R}+m$, where $L_{R}$ is the level of $Q(x,R)$. Since $|Q(x,R)|\geq 3^{-1}c_{0}\delta^{L_{R}}$ by Theorem A (iii) and (2.1), we have by (2.4) and another application of Theorem A (iii) on $Q_{i}$ that
@@ 4. Final Remarks
The construction of dyadic cube systems of Hytönen-Kairema in [24] is actually given for quasimetric spaces. As a result, Theorem 1.1 is also true if $X$ is a quasimetric doubling space. The proofs are almost identical, with the only difference being the dependence of a few of the uniform constants on the quasimetric constant of the space. Another generalization of Theorem 1.1 would be to state it for all $\delta\in(0,1)$ small enough that satisfy the relation $12C_{0}\delta\leq c_{0}$ in Theorem A, which follows with identical arguments to those in Section 3.